A rectangle with a perimeter 38 cm is 3 cm longer than it is wide. Determine the dimensions of the rectangle.
Answer by mathsmiles(68)

P of a rectangle is the L + L + W + W or 2L + 2W

Here we know that P is 38.
We also know that L is 3cm+W
L = 3+W

So substituting in the formula for perimeter:
2L + 2W = P = 38
2(3+W) + 2W = 38 Now multiply out the paren
6 + 2W + 2W = 38 Combine W terms
6 + 4W = 38
Subtracting 6 from both sides of the equation:
4W = 32
Dividing both sides by 4:
W = 8cm

But we know the length is W + 3cm
L = 8 + 3 = 11cm

Checking:
2(8) + 2(11) = 38
16 + 22 = 38
38 = 38 Correct!